OpenStudy (anonymous):

Simplify the rational expression. State any restrictions on the variable. n^4 - 10n^2 + 24 / n^4 - 9n^2 + 18

Parth (parthkohli):

You have to factor the numerator and denominator. For that, you can do algebraic substitution. Start with the numerator: Let \(n^2 = y\), then we have \(y^2 - 10y + 24\)... factor this quadratic expression. Same for the denominator: Let \(n^2 = y\), then we have \(y^2 - 9n + 18\)

Parth (parthkohli):

For the restrictions, the denominator can't be zero, so the restriction is the following:\[y^2 - 9n + 18 = 0\]Solve for \(y\), then square it to solve for \(n^2\).

Parth (parthkohli):

Actually, find the square root of \(y\) to solve for \(n\).

OpenStudy (anonymous):

\[\frac{ (y-6)(y-4)}{ (y-3)(y-6) }\]